Operations Report – May 11th

Crew 212 Operations Report 11-MAY-2019

SOL: 6

Name of person filing report: Paolo Guardabasso

Non-nominal systems: None

Notes on non-nominal systems: Nothing to report

Generator (hours run last night): 9:30

On (last night): 9:00 PM

Off (this morning): 6:30 AM

Solar— SOC 88% at 7:30 PM

Diesel Reading – 95% [From Director]

Propane Reading – 45%

Ethanol Free Gasoline – 4 gallons

Water (Axillary tank) – 0 gallons

Water (static tank) – 170 gallons

Axillary to Static tank transfer – no

Gallons transferred: No

Water in GreenHab – 150 gallons

Water (loft) – Static to Loft Pump used – Yes

Water Meter: 143500.4

Toilet tank emptied: Yes

Deimos rover used: no (out of order)


Beginning charge:

Ending charge:

Currently charging:

Sojourner rover used: ASSIGNED TO DIRECTOR


Beginning charge:

Ending charge:

Currently charging:

Spirit rover used: Yes

Hours: 92.8h

Beginning charge: 100%

Ending charge: 79%

Currently charging: Yes

Opportunity rover used: Yes

Hours: 59.1h

Beginning charge:

Ending charge:

Currently charging: Yes

Curiosity rover used: Yes


Beginning charge:

Ending charge:

Currently charging: Yes

Notes on rovers: Curiosity and Opportunity were missing one wheel hubcap per rover. It’s not yet clear exactly how they went missing but the Director is dealing with the matter and the one on Curi has already been replaced. A new charging cable was added in order to be able to charge 3 rovers at the same time.

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3)

Reason for use:

Oil Added?

ATV Fuel Used: Gals

# Hours the ATVs were Used today:

Notes on ATVs:

HabCar used and why, where? Nothing to report

CrewCar used and why, where? Off-Campus

General notes and comments: The crew worked on their experiment and an EVA was delayed and lasted 1 hour instead of 2.

Summary of internet: Nothing to report.

Summary of suits and radios: Some radios didn’t hold their charge, major testing will be carried out in the coming days. Radios proved not working at a range of about 1 km from the Hab during EVA. Some spare batteries for suits were not able to recharge.

Summary of Hab operations: Fixing of the oven door is still ongoing.

Summary of GreenHab operations: The fan was turned on as the temperature passed 40° C and it managed to cool the room to 30°C.

Summary of ScienceDome operations: Microbiological experiments were run.

Summary of RAM operations: Nothing to report.

Summary of any observatory issues: The crew astronomer reported that the telescope is not properly focused.

Summary of health and safety issues: Nothing to report.

Questions, concerns, and requests to Mission Support: Several tourist cars approached the MDRS; the one that passed the sign was stopped by the Director on its way to the Hab.

[status draft

Today was Sol 6. We started the day with pancakes and since we are all feeling weak from the reduced nutrients we tried an attempt at banana smoothies.

In the solar telescope a photograph was taken of the sun however it was out of focus and another attempt will be made after suggestions from Peter. In addition, 65 Pelican nebula pictures have been processed awaiting the final 20 photos tonight.

Each crew member worked on their own projects today, the drones were given minor repairs from yesterday’s EVAs and batteries charged from today’s. The Martian stimulant soil has been prepared and sterilized, tomato and two different kinds of lettuce have been sowed with the soil and bacteria.

A group helped to reinstall the software for the Rover, progress is being made with the potential of success in the next few days.

Today one EVA took place, in an attempt to break the ice in the pressurization time the crew told “Dad jokes” to the EVA team. Jokes included: Why did the skeleton not go to the party? Because he has nobody to go with. These classics were not appreciated by the rest of the team! The EVA was to test the first autonomous mission, this was taken near Kissing camel and was a success.

Once returned we completed another round of the problem-solving game.

Kind Regards,
Zoe Townsend

Science Report – May 11th

Crew 212-LATAM-III

Submitted by Spacesuit Engineer: Camilo Andrés Zorro Mendoza

Spacesuit report

Sol: 6

Sim suit used: 3 and 5.
used for 1 hour

Sim suit: 3
Voltage level before EVA: 13.6 V
Voltage level after EVA: 12.9 V

Sim suit: 5
Voltage level before EVA: 13. V
Voltage level after EVA: 12.8 V

General commentaries

The batteries of radios are not working properly. In the following days, I will give you a detail report on the number of batteries that should be changed.

Greenhab Report – May 11th

Crew 212 GreenHab Report – 11-May-2019


GreenHab Officer: Hector Palomeque

Environmental control: Ambient with the door open + fan turned up to 3 (max) from 14:00 to 17:30

Shade cloth 80%: on

Average temperatures:

Low: 23.5 °C (±0.45)

High: 41.3 °C (±1.17)

Hours of supplemental light: 0

Daily water usage for crops: 4 gallons + 2 gallons

Water in Blue Tank: 151 gallons

Time(s) of watering for crops: 11:30 + 19:00

Changes to crops: All plants look healthy. Basil and tomato plants continue blooming. The 2 tomato plants in the black-metallic shelf are about to ripen its fruits (+10 tomatoes are yellow). Tomato in the big blue tray seems to be recovered from the hydric stress.

Narrative: The GreenHab fan had to be turned on because of the indoor temperature has reached + 40 ° C. One hour after turning the fan on, the indoor temperature decreased to about +30°C.

Plants were watered half hour before noon and those that were dry at the end of the day (7:00 PM) were watered again. The usual watering at the early morning (at 07:00) was avoided because of the substrate has remained moisture from the day before.

Chard harvested for preparing the lunch of the Crew.

Harvested: 15g Chard

Support/supplies needed: None.

I have attached an excel sheet with today’s temperature tracking of the GreenHab.


Sol Summary – May 11th

Crew 212 Sol 6 Summary Report 10-MAY-2019

Sol: 6

Summary Title: Finally, some nice sky

Author’s name: Marlen Castillo Vilcahuaman

Mission Status: A-ok

Sol Activity Summary: For today, we had an EVA in the afternoon. We decided to experiment with some new recipes in the kitchen. We’re glad there’s some sun after some cloudy days.

Look Ahead Plan: An EVA for exploring around the hab some other plant sites.

Anomalies in work: Today, we found that there were 2 missing parts from the Rovers (see Operations Report). This was communicated by the Director. One part was found. The director is currently searching for the other one.

Weather: Sun is shining up.

Crew Physical Status: OK

EVA: 1

Reports to be file: Operations Report, GreenHab Report, Journalist Report, Daily Photo Report, EVA Request, Space Suit Report, EVA Report, Astronomy Report

Support Requested: We have 1 bag of burnable trash and 1 bag of non-burnable trash.

Astronomy Report – May 11th

Astronomy Report
Name: Mariona Badenas
Crew: 212
Date: 11 May 2019

Robotic Telescope Requested (choose one): MDRS-WF

Object Viewed #1: IC5067

Problems Encountered: The telescope was open at the start of the night but did not take pictures of this nebula (this problem has already been discussed with Peter).

(More objects to be Viewed): NGC6992, Veil Nebula, Iris Nebula
Problems Encountered: N/A (Observations of these objects have not begun yet).

Solar Features Observed: Observed the sun through the eyepiece and took several pictures (see attached photo) but did not succeed at properly focusing the camera.

Problems Encountered: Camera wasn’t focused well. Will attempt to fix this issue tomorrow following Peter’s instructions.

Copyright © The Mars Society. All rights reserved. | Main Site